A 2.0-m length of wire is made by welding the end of a 120-cm-long silver wire to the end of an 80-cm-long copper wire. Each piece of wire is 0.60 mm in diameter. The wire is at room temperature, so the resistivities are as given in Table 25.1. A potential difference of 9.0 V is maintained between the ends of the 2.0-m composite wire.
Answer #1

Resistivity of silver wire = 1.47e-8 ohm.meter

Resistivity of copper wire = 1.72e-8 ohm.meter

we know that

Resistance = resisitivity * length / cross sectional area

R = L/A

so,

I = A ( V / AgLAg + CuLCu)

I = 81.04 A

Therefore,

current through copper and silver wire will be 81.04 A ( same because they are in series)

------------------------------------------------------------------------------------------------------------------------------

For E calculation

For copper

E = resistivity * current density

E = resistivity * current / area

E = 4.9298 V/m

For silver

E = 4.213 V/m

------------------------------------------------------------------------------

Potential difference = E * L

for silver

Potential difference = 4.213 * 1.2 = 5.055 V

For copper

Potential difference = 3.9438 V